# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v5

@@ Abstract
Language models (LMs) risk inadvertently memorizing and divulging sensitive or personally identifiable information (PII) seen in training data, causing privacy concerns. Current approaches to address this issue involve costly dataset scrubbing, or model filtering through unlearning and model editing, which can be bypassed through extraction attacks. We propose REVS, a novel non-gradient-based method for unlearning sensitive information from LMs. REVS identifies and modifies a small subset of neurons relevant for constituent tokens that form sensitive information. To adequately evaluate our method on truly sensitive information, we curate three datasets: email and URL datasets naturally memorized by the models, and a synthetic social security number dataset that we tune the models to memorize. Compared to other methods, REVS demonstrates superior performance in unlearning sensitive information and robustness to extraction attacks, while retaining underlying model integrity.
@@ 1 Introduction
We introduce Rank Editing in the Vocabulary Space (REVS), a novel non-gradient-based model-editing approach that enables robust unlearning of sensitive information from LMs while maintaining model performance and offering strong robustness against extraction attacks. Adopting the view that transformer MLP layers construct predictions by promoting specific tokens in the output vocabulary space (Geva et al., 2022), REVS locates the layers and particular subsets of neurons that promote the tokens corresponding to the targeted sensitive information. By modifying these neurons with respect to the relevant tokens, REVS surgically removes the model’s encoded tendency to generate that sensitive data while preserving its broader knowledge and remaining robust to extraction attacks. See Figure 1 for an illustration.
@@ 2 Problem Setup and Preliminaries
Consider the case where a LM is prompted with text from its training dataset that contains sensitive information, e.g., “Contact David Lewis at ”. If the model has memorized the completion of this prompt, it will generate the relevant email address, such as “lewis.david@email.com”, illustrated in Figure 1. This scenario highlights a prevalent issue with LMs: their ability to memorize can lead to unintended divulgence of sensitive data, in particular PII. To address this issue, dataset scrubbing approaches tackle it from a data perspective, preventing sensitive information from being stored in the model during training. Machine unlearning approaches take on a post-hoc approach, modifying either all, or just a subset of parameters that contain sensitive information. The goal of these approaches is to efficiently and effectively prevent the model from generating each sequence containing sensitive information in a way robust to extraction attacks, while preserving general capabilities of the model (Liu et al., 2024).
@@
We briefly describe the main components in Transformer LMs; more details are found elsewhere (Vaswani et al., 2017; Black et al., 2021). Omitting some technical details, a Transformer LM is made of a sequence of blocks. Each block contains self-attention and multi-layer perceptrons (MLPs) that write and read to the residual stream (Elhage et al., 2021). The MLP is defined as $\text{MLP}(\vec{x})$ $=\text{FF}_{2}\sigma(\text{FF}_{1}\vec{x})$, and is added to the residual stream. The final layer hidden state is projected to the vocabulary space via the unembedding matrix $U$, followed by a softmax to obtain next token probabilities.
@@ 3 Methodology
The higher the rank, the more likely the token is to be generated, where rank $r_{t}=1$ denotes the most likely token. To unlearn sensitive information, REVS aims to increase $r_{t}$ value to a desired $r_{d}$, thus decreasing its rank by identifying and modifying $FF_{2}$ columns that contribute most to generating the target token $t$. This unlearning process operates in two main phases: localization (Section 3.1), where we identify the layers and neurons ($FF_{2}$ columns) that contain information relevant for generating $t$, and editing (Section 3.2), where we overwrite information relevant for $t$ while not affecting knowledge used for generating other tokens. We provide a high-level overview of REVS in Algorithm 1.
@@
8:              $\vec{n}\leftarrow\mathrm{SelectNeuron}(M,P,l,t)$triangleright Sec.3.1.2
9:              $\vec{n}\leftarrow\mathrm{EditNeuron}(\vec{n},t)$triangleright Sec.3.2
10:              $r_{t}\leftarrow\mathrm{TokenRank}(M,P,l,t)$triangleright Eq.1
@@ 3.1.2 Neuron Selection
This hybrid selection approach targets neurons that are both contextually significant and semantically relevant, thus outperforming alternatives such as selection based solely on activations, token associations, gradients (Dai et al., 2022), or random selection (see Section D.1).
@@ 3.2 Neuron Editing
For each selected neuron $\vec{n}_{j}$, REVS performs iterative adjustment as illustrated in Figure 2:
@@
We initialize $l_{t}$ with a small logit value corresponding to a large $r_{t}$ value (low rank), and iteratively decrease it by multiplying with a constant factor in each iteration until $r_{t}>r_{n}$, where $r_{n}$ is the neuron’s desired rank. For a detailed description of the iterative process see Section B.2.
@@ 3.3 Target Token Selection
Sensitive information typically spans multiple tokens, but unlearning every token is unnecessary, as accurately recovering the original data requires reconstructing the full token sequence. To account for this, for each target sequence $S$, we select a subset of the $T\subseteq S$ rarest tokens, as unlearning targets for all considered methods, with $|T|=2$ in this work. See Sections E.1 and E.2 for details and ablation experiments.
@@ 4.2 Datasets
Organically memorized data: Identifying naturally memorized sensitive information requires two computationally intensive steps: detecting genuine sensitive information and verifying its memorization. Due to computational constraints, we use a pre-filtered subset of The Pile (Gao et al., 2020) containing 15,000 sentences.111github.com/google-research/lm-extraction-benchmark From this, we identified memorized instances in Llama-3-8B (205 email addresses, 203 URLs) and GPT-J-6B (288 email addresses). Synthetic SSN dataset: We generated 200 sentences containing social security numbers using Claude 3 Sonnet (Anthropic, 2024) and fine-tuned the base models to memorize them. See Appendix F for more details.
@@ 4.3 Baselines
We elaborate on the used baselines and their hyperparameters in Appendix C.
@@ 4.4.1 Unlearning Effectiveness
[…]
@@
[…]
Generalization: We evaluate unlearning generalization by measuring Efficacy@$k$ on the target token set $T$ using prompts that were unseen during the unlearning process but are tuned to generate the same sensitive information (see Section F.2, Table 12 for examples).
@@ 4.4.3 Extraction Resistance
SSN Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $100{\pm 0.00}$ 61.0561.0561.05{}61.05 $47.83$
FT-L $\underline{36.98{\pm 11.97}}$ $\underline{63.88{\pm 9.88}}$ $\underline{50.35{\pm 10.76}}$ $24.33{\pm 9.78}$ 60.9960.9960.99{}60.99 $46.62$
MEMIT $24.72{\pm 7.21}$ $30.70{\pm 9.67}$ $23.90{\pm 7.61}$ $22.67{\pm 6.50}$ 61.0261.0261.02{}61.02 $46.17$
Max-Entropy $5.12{\pm 2.13}$ $5.17{\pm 3.00}$ $3.92{\pm 2.33}$ $1.40{\pm 0.60}$ 61.0661.0661.06{}61.06 $47.46$
Head-Projection $2.98{\pm 0.79}$ $3.08{\pm 1.23}$ $2.95{\pm 0.68}$ $4.17{\pm 2.41}$ 61.0661.0661.06{}61.06 $46.92$
RMU $16.42{\pm 9.10}$ $13.47{\pm 8.42}$ $16.67{\pm 10.41}$ $\underline{38.67{\pm 14.92}}$ 60.8360.8360.83{}60.83 $48.21$
NPO-KL $11.95{\pm 4.87}$ $38.78{\pm 18.34}$ $36.13{\pm 16.59}$ $6.33{\pm 3.68}$ 61.0161.0161.01{}61.01 $47.23$
REVS (ours) $\bm{89.58^{\dagger}{\pm 1.99}}$ $\bm{98.88{\pm 1.28}}$ $\bm{89.67{\pm 3.78}}$ $\bm{82.17{\pm 5.08}}$ 60.8760.8760.87{}60.87 $44.20$
Emails Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$- $\bm{100{\pm 0.00}}$ 62.1762.1762.17{}62.17 47.9947.9947.99{}47.99
FT-L $\underline{50.30{\pm 3.04}}$ $52.98{\pm 4.23}$- ${49.25{\pm 8.50}}$ 62.1562.1562.15{}62.15 $50.94$
MEMIT $35.43{\pm 4.30}$ $63.63{\pm 3.50}$- $24.84{\pm 4.20}$ 62.2262.2262.22{}62.22 $50.64$
Max-Entropy $31.08{\pm 3.30}$ $\bm{69.75{\pm 6.30}}$- $20.22{\pm 3.10}$ 62.1162.1162.11{}62.11 $50.64$
Head-Projection $30.80{\pm 3.90}$ $\underline{64.33{\pm 4.90}}$- $20.43{\pm 3.40}$ 62.1062.1062.10{}62.10 $50.19$
RMU $17.47{\pm 3.60}$ $15.08{\pm 5.90}$- $32.58{\pm 16.00}$ 62.1062.1062.10{}62.10 $46.39$
NPO-KL $32.75{\pm 2.70}$ $24.27{\pm 3.00}$- $\underline{50.97{\pm 2.00}}$ 62.0562.0562.05{}62.05 $48.67$
REVS (ours) $\bm{62.37^{\dagger}{\pm 2.30}}$ $59.65{\pm 3.95}$- $65.70{\pm 3.79}$ 61.7761.7761.77{}61.77 47.4647.4647.46{}47.46
@@
REVS (ours) $\bm{44.25^{\dagger}{\pm 5.01}}$ $\bm{78.22{\pm 6.04}}$- $30.94{\pm 4.11}$ $62.31$ 47.7647.7647.76{}47.76
@@ 4.5 Implementation Details
In each experiment, we unlearn all sensitive sequences before evaluation. The Unlearning Score is the harmonic mean of efficacy, specificity, and generality (for the SSN dataset), while the Resistance Score is the harmonic mean of all extraction attacks. Hyperparameters (HPs) are optimized per method on the SSN and Email datasets to maximize Unlearning Score with $k=100$, using a fixed seed (00). In the URL dataset, we use the same HPs as the Emails dataset, to assess robustness and generalization across naturally memorized PII types. We cross-validate over 6 random splits (seeds 11116666). Each split has a forget set of $50$ targets in Emails and URLs, while in SSN, it comprises half the data. The forget set is used for efficacy, generalization (SSN), and extraction resistance, while the rest is used for specificity.
@@ 5 Results
Figure 3: Results for Extraction Resistance on Llama-3-8B for $k=100$. REVS is more robust to extraction attacks. We report the full results in tabular format in Table 2 (Section A.2).
@@
As Table 1 and Figure 3 show, REVS outperforms all baselines in Unlearning Score and is on par or better in Resistance Score, for Llama-3-8B across all datasets. Results on GPT-J-6B follow the same trends (Section A.3).
@@
REVS achieves the best efficacy and generality on SSN and URLs and strong efficacy on Emails. As shown in Figure 4, perfect efficacy is possible but compromises specificity. Notably, methods surpassing REVS in efficacy on Emails achieve a lower Unlearning Score due to poor specificity.
@@
As shown in Figure 3 and Table 2, REVS achieves the highest Resistance Score across all datasets and models, except for Emails on Llama-3-8B, where it ranks a close second. Notably, Tables 2 and 4 (Section A.2) highlight a clear link between higher Effectiveness and improved Resistance Score. While strong Effectiveness often compromises specificity, REVS maintains both, achieving the best or second-best Resistance Score and demonstrating balanced robust unlearning.
@@
Samples of generated text after unlearning are in Section B.3, showing the model generates coherent alternative text instead of memorized private emails.
@@ 5.1 Analysis
In this section we compare the robustness of REVS with FT-L, the second-best performing method, under different hyperparameters. Figure 4 shows that REVS Pareto-dominates FT-L in all but one experiment.
@@
REVS consistently outperforms FT-L as we increase the size of the candidate token set $C$ (Figures 4, 4 and 4). While for the Emails and URLs datasets both methods exhibit similar trends, REVS maintains superior performance for each size of $C$. For the SSN dataset, REVS demonstrates near-perfect efficacy across all candidate token set sizes.
@@
While both REVS and FT-L are affected by this trade-off (Figures 4, 4 and 4), REVS fares better, robustly maintaining high specificity when increasing edit strength. Notably, on SSN, there is almost no trade-off, with REVS achieving near-perfect specificity and efficacy.
@@
On SSN (Figure 4) REVS maintains a stable Unlearning Score across edit counts, while FT-L performs well only at its tuned hyperparameters, degrading elsewhere. Organic data (Figures 4 and 4) show a downward trend in Unlearning Score for both methods as edits increase; however, REVS consistently outperforms FT-L.

### v6

@@ Abstract
Language models (LMs) risk inadvertently memorizing and divulging sensitive or personally identifiable information (PII) seen in training data, causing privacy concerns. Current approaches to address this issue involve costly dataset scrubbing, or model filtering through unlearning and model editing, which can be bypassed through extraction attacks. We propose REVS, a novel non-gradient-based method for unlearning sensitive information from LMs. REVS identifies and modifies a small subset of neurons relevant for constituent tokens that form sensitive information. To adequately evaluate our method on truly sensitive information, we curate three datasets: email and URL datasets naturally memorized by the models, and a synthetic social security number dataset that we tune the models to memorize. Compared to other methods, REVS demonstrates superior performance in unlearning sensitive information and robustness to extraction attacks, while retaining underlying model integrity. [^1] Code available at github.com/tomerashuach/REVS.
@@ 1 Introduction
We introduce Rank Editing in the Vocabulary Space (REVS), a novel non-gradient-based model-editing approach that enables robust unlearning of sensitive information from LMs while maintaining model performance and offering strong robustness against extraction attacks. Adopting the view that transformer MLP layers construct predictions by promoting specific tokens in the output vocabulary space (Geva et al., 2022), REVS locates the layers and particular subsets of neurons that promote the tokens corresponding to the targeted sensitive information. By modifying these neurons with respect to the relevant tokens, REVS surgically removes the model’s encoded tendency to generate that sensitive data while preserving its broader knowledge and remaining robust to extraction attacks. See Figure˜1 for an illustration.
@@ 2 Problem Setup and Preliminaries
Consider the case where a LM is prompted with text from its training dataset that contains sensitive information, e.g., “Contact David Lewis at ”. If the model has memorized the completion of this prompt, it will generate the relevant email address, such as “lewis.david@email.com”, illustrated in Figure˜1. This scenario highlights a prevalent issue with LMs: their ability to memorize can lead to unintended divulgence of sensitive data, in particular PII. To address this issue, dataset scrubbing approaches tackle it from a data perspective, preventing sensitive information from being stored in the model during training. Machine unlearning approaches take on a post-hoc approach, modifying either all, or just a subset of parameters that contain sensitive information. The goal of these approaches is to efficiently and effectively prevent the model from generating each sequence containing sensitive information in a way robust to extraction attacks, while preserving general capabilities of the model (Liu et al., 2024).
@@
We briefly describe the main components in Transformer LMs; more details are found elsewhere (Vaswani et al., 2017; Black et al., 2021). Omitting some technical details, a Transformer LM is made of a sequence of blocks. Each block contains self-attention and multi-layer perceptrons (MLPs) that write and read to the residual stream (Elhage et al., 2021). The MLP is defined as $\text{MLP}(\vec{x})$ $\text{FF}_{2}\sigma(\text{FF}_{1}\vec{x})$, and is added to the residual stream. The final layer hidden state is projected to the vocabulary space via the unembedding matrix $U$, followed by a softmax to obtain next token probabilities.
@@ 3 Methodology
The higher the rank, the more likely the token is to be generated, where rank $r_{t}=1$ denotes the most likely token. To unlearn sensitive information, REVS aims to increase $r_{t}$ value to a desired $r_{d}$, thus decreasing its rank by identifying and modifying $FF_{2}$ columns that contribute most to generating the target token $t$. This unlearning process operates in two main phases: localization (Section˜3.1), where we identify the layers and neurons ($FF_{2}$ columns) that contain information relevant for generating $t$, and editing (Section˜3.2), where we overwrite information relevant for $t$ while not affecting knowledge used for generating other tokens. We provide a high-level overview of REVS in Algorithm˜1.
@@
8:     $\vec{n}\leftarrow\mathrm{SelectNeuron}(M,P,l,t)$ $\triangleright$ Sec.3.1.2
9:     $\vec{n}\leftarrow\mathrm{EditNeuron}(\vec{n},t)$ $\triangleright$ Sec.3.2
10:     $r_{t}\leftarrow\mathrm{TokenRank}(M,P,l,t)$ $\triangleright$ Eq.1
@@ 3.1.2 Neuron Selection
This hybrid selection approach targets neurons that are both contextually significant and semantically relevant, thus outperforming alternatives such as selection based solely on activations, token associations, gradients (Dai et al., 2022), or random selection (see Section˜D.1).
@@ 3.2 Neuron Editing
For each selected neuron $\vec{n}_{j}$, REVS performs iterative adjustment as illustrated in Figure˜2:
@@
We initialize $l_{t}$ with a small logit value corresponding to a large $r_{t}$ value (low rank), and iteratively decrease it by multiplying with a constant factor in each iteration until $r_{t}>r_{n}$, where $r_{n}$ is the neuron’s desired rank. For a detailed description of the iterative process see Section˜B.2.
@@ 3.3 Target Token Selection
Sensitive information typically spans multiple tokens, but unlearning every token is unnecessary, as accurately recovering the original data requires reconstructing the full token sequence. To account for this, for each target sequence $S$, we select a subset of the $T\subseteq S$ rarest tokens, as unlearning targets for all considered methods, with $|T|=2$ in this work. See Sections˜E.1 and E.2 for details and ablation experiments.
@@ 4.2 Datasets
Organically memorized data: Identifying naturally memorized sensitive information requires two computationally intensive steps: detecting genuine sensitive information and verifying its memorization. Due to computational constraints, we use a pre-filtered subset of The Pile (Gao et al., 2020) containing 15,000 sentences.222github.com/google-research/lm-extraction-benchmark From this, we identified memorized instances in Llama-3-8B (205 email addresses, 203 URLs) and GPT-J-6B (288 email addresses). Synthetic SSN dataset: We generated 200 sentences containing social security numbers using Claude 3 Sonnet (Anthropic, 2024) and fine-tuned the base models to memorize them. See Appendix˜F for more details.
@@ 4.3 Baselines
We elaborate on the used baselines and their hyperparameters in Appendix˜C.
@@ 4.4.1 Unlearning Effectiveness
$\text{Score@$k$}(t)=\begin{cases}r_{t}/k,&\text{if }k>r_{t}\\ 1,&\text{otherwise}\end{cases}$ (2)
@@
$\textrm{Efficacy@$k$}=\max_{t\in T}\text{Score}@k(t)$ (3)
Generalization: We evaluate unlearning generalization by measuring Efficacy@$k$ on the target token set $T$ using prompts that were unseen during the unlearning process but are tuned to generate the same sensitive information (see Section˜F.2, Table˜12 for examples).
@@ 4.4.3 Extraction Resistance
SSN Unedited ${\pm 0.00} 0.00±0.00$$\pm 0.00$} $0.00{\pm 0.00}$ $100{\pm 0.00}$ 61.0561.05{} $47.83$
FT-L $\underline{36.98{\pm 11.97}}$ $\underline{63.88{\pm 9.88}}$ $\underline{50.35{\pm 10.76}}$ $24.33{\pm 9.78}$ 60.9960.99{} $46.62$
MEMIT $24.72{\pm 7.21}$ $30.70{\pm 9.67}$ $3.90{\pm 7.61} 22$$2.67{\pm 6.50}$ 61.0261.02{} $46.17$
Max-Entropy $5.12{\pm 2.13}$ $5.17{\pm 3.00}$ $3.92{\pm 2.33}$ $1.40{\pm 0.60}$ 61.0661.06{} $47.46$
Head-Projection $2.98{\pm 0.79}$ $3.08{\pm 1.23}$ $2.95{\pm 0.68}$ $4.17{\pm 2.41}$ 61.0661.06{} $46.92$
RMU $16.42{\pm 9.10}$ $13.47{\pm 8.42}$ $16.67{\pm 10.41}$ $\underline{38.67{\pm 14.92}}$ 60.8360.83{} $48.21$
NPO-KL $11.95{\pm 4.87}$ $38.78{\pm 18.34}$ $36.13{\pm 16.59}$ $6.33{\pm 3.68}$ 61.0161.01{} $47.23$
REVS (ours) $\bm{89.58^{\dagger}{\pm 1.99}}$ $\bm{98.88{\pm 1.28}}$ $\bm{89.67{\pm 3.78}}$ $\bm{82.17{\pm 5.08}}$ 60.8760.87{} $44.20$
Emails Unedited ${\pm 0.00} 0.00±0.00$$\pm 0.00$} - $\bm{100{\pm 0.00}}$ 62.1762.17{} 47.9947.99{}
FT-L $\underline{50.30{\pm 3.04}}$ $52.98{\pm 4.23}$ - $49.25{\pm 8.50}$} 62.1562.15{} $50.94$
MEMIT $35.43{\pm 4.30}$ $63.63{\pm 3.50}$ - $24.84{\pm 4.20}$ 62.2262.22{} $50.64$
Max-Entropy $31.08{\pm 3.30}$ $\bm{69.75{\pm 6.30}}$ - $20.22{\pm 3.10}$ 62.1162.11{} $50.64$
Head-Projection $30.80{\pm 3.90}$ $\underline{64.33{\pm 4.90}}$ - $20.43{\pm 3.40}$ 62.1062.10{} $50.19$
RMU $17.47{\pm 3.60}$ $15.08{\pm 5.90}$ - $32.58{\pm 16.00}$ 62.1062.10{} $46.39$
NPO-KL $32.75{\pm 2.70}$ $24.27{\pm 3.00}$ - $\underline{50.97{\pm 2.00}}$ 62.0562.05{} $48.67$
REVS (ours) $\bm{62.37^{\dagger}{\pm 2.30}}$ $59.65{\pm 3.95}$ - $65.70{\pm 3.79}$ 61.7761.77{} 47.4647.46{}
@@
REVS (ours) $\bm{44.25^{\dagger}{\pm 5.01}}$ $\bm{78.22{\pm 6.04}}$ - $30.94{\pm 4.11}$ $62.31$ 47.7647.76{}
@@ 4.5 Implementation Details
In each experiment, we unlearn all sensitive sequences before evaluation. The Unlearning Score is the harmonic mean of efficacy, specificity, and generality (for the SSN dataset), while the Resistance Score is the harmonic mean of all extraction attacks. Hyperparameters (HPs) are optimized per method on the SSN and Email datasets to maximize Unlearning Score with $k=100$, using a fixed seed (0). In the URL dataset, we use the same HPs as the Emails dataset, to assess robustness and generalization across naturally memorized PII types. We cross-validate over 6 random splits (seeds 1166). Each split has a forget set of $50$ targets in Emails and URLs, while in SSN, it comprises half the data. The forget set is used for efficacy, generalization (SSN), and extraction resistance, while the rest is used for specificity.
@@ 5 Results
Figure 3: Results for Extraction Resistance on Llama-3-8B for $k=100$. REVS is more robust to extraction attacks. We report the full results in tabular format in Table˜2 (Section˜A.2).
@@
As Table˜1 and Figure˜3 show, REVS outperforms all baselines in Unlearning Score and is on par or better in Resistance Score, for Llama-3-8B across all datasets. Results on GPT-J-6B follow the same trends (Section˜A.3).
@@
REVS achieves the best efficacy and generality on SSN and URLs and strong efficacy on Emails. As shown in Figure˜4, perfect efficacy is possible but compromises specificity. Notably, methods surpassing REVS in efficacy on Emails achieve a lower Unlearning Score due to poor specificity.
@@
As shown in Figure˜3 and Table˜2, REVS achieves the highest Resistance Score across all datasets and models, except for Emails on Llama-3-8B, where it ranks a close second. Notably, Tables˜2 and 4 (Section˜A.2) highlight a clear link between higher Effectiveness and improved Resistance Score. While strong Effectiveness often compromises specificity, REVS maintains both, achieving the best or second-best Resistance Score and demonstrating balanced robust unlearning.
@@
Samples of generated text after unlearning are in Section˜B.3, showing the model generates coherent alternative text instead of memorized private emails.
@@ 5.1 Analysis
In this section we compare the robustness of REVS with FT-L, the second-best performing method, under different hyperparameters. Figure˜4 shows that REVS Pareto-dominates FT-L in all but one experiment.
@@
REVS consistently outperforms FT-L as we increase the size of the candidate token set $C$ (Figures˜4, 4 and 4). While for the Emails and URLs datasets both methods exhibit similar trends, REVS maintains superior performance for each size of $C$. For the SSN dataset, REVS demonstrates near-perfect efficacy across all candidate token set sizes.
@@
While both REVS and FT-L are affected by this trade-off (Figures˜4, 4 and 4), REVS fares better, robustly maintaining high specificity when increasing edit strength. Notably, on SSN, there is almost no trade-off, with REVS achieving near-perfect specificity and efficacy.
@@
On SSN (Figure˜4) REVS maintains a stable Unlearning Score across edit counts, while FT-L performs well only at its tuned hyperparameters, degrading elsewhere. Organic data (Figures˜4 and 4) show a downward trend in Unlearning Score for both methods as edits increase; however, REVS consistently outperforms FT-L.